OpenStudy (anonymous):

The determinant of the matrix (p 10) is equal to 4 (2 3)

OpenStudy (anonymous):

the value of p is:

OpenStudy (anonymous):

only square matrix have determinant

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

\[(p \times 3) - (10\times 2) = 4\]

OpenStudy (anonymous):

thank you, i have got it. the answer is 8?

OpenStudy (anonymous):

yup